OpenStudy (anonymous):

What are the points of intersection of the graphs of the two functions? f(x)=g(x) f(x)=x^2+6x+1 g(x)=1 I set them equal and then got x^2+6x=0, but where do I go from there? Help please. = [

OpenStudy (unklerhaukus):

good start\[x^2+6x=0\]\[x^2 =-6x\]\[x=-6,~~~~-x=-6\]\[~~~~~~~~~~~~~~~~~~~~~~~x=6\] \[x=±6 \]

OpenStudy (anonymous):

oh!!! thanks so much!! = ]

OpenStudy (unklerhaukus):

actally i think i have made a mistake

OpenStudy (unklerhaukus):

i cant see where it is though

OpenStudy (anonymous):

x=-6 and x=0. x^2+6x=0 x(x+6)=0 x=0 and x=-6

OpenStudy (unklerhaukus):

shankvee has the right answer

OpenStudy (anonymous):

oh i see... thanks shankvee

OpenStudy (unklerhaukus):

i still cant see what my mistake was

OpenStudy (anonymous):

I think it was because it was a x^2 instead of x just being a variable to a constant...maybe? idk. Wouldn't the left side of the equation would of had to of been something like (2x+4)^2 in order for the solving using the square root?

OpenStudy (unklerhaukus):

i wasnt taking square root thoug, i was dividing by x

OpenStudy (anonymous):

x^2+6x =0 x( x + 6) = 0 x = 0, x = -6

OpenStudy (anonymous):

so when it asks for the intersection, what do i plug in for g(x)=1 since the x-intercepts are 0,-6? just multiply?

OpenStudy (anonymous):

nevermind, the y intercept would just be 1

OpenStudy (anonymous):

It intersects when x=0 and x=-6. If you want the points they are (0,1) and (-6,1).

OpenStudy (anonymous):

thanks!